OpenStudy (anonymous):

Help please! find the range of function sqrt(9-x^2) on interval -3≤x ≤2

OpenStudy (anonymous):

\(y=\sqrt{9-x^2}\) is the top half of a circle with radius 3 centered at the origin. Do you see why the range of the circle is \(0\le y\le3\) over the interval \(-3\le x\le3\) ?

OpenStudy (anonymous):

 \(y\) doesn't get higher than 3.

OpenStudy (anonymous):

So the range you said is the same for the interval -3≤x ≤2?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw